# Narrative Scaffolding: Transforming Data-Driven Sensemaking Through Narrative-First Exploration

### v1

@@ Abstract.
When exploring data, analysts construct narratives about what the data means by asking questions, generating visualizations, reflecting on patterns, and revising their interpretations as new insights emerge. Yet existing analysis tools treat narrative as an afterthought, breaking the link between reasoning, reflection, and the evolving story from exploration. Consequently, analysts lose the ability to see how their reasoning evolves, making it harder to reflect systematically or build coherent explanations. To address this gap, we propose Narrative Scaffolding, a framework for narrative-driven exploration that positions narrative construction as the primary interface for exploration and reasoning. We implement this framework in a system that externalizes iterative reasoning through narrative-first entry, semantically aligned view generation, and reflection support via insight provenance and inquiry tracking. In a within-subject study ($N=20$), we demonstrate that narrative scaffolding facilitates broader exploration, deeper reflection, and more defensible narratives. An evaluation with visualization literacy experts (N = 6) confirmed that the system produced outputs aligned with narrative intent and facilitated intentional exploration.
@@ 1. Introduction
Exploring data is fundamentally a sensemaking process that involves practicing interpretive reasoning throughout: in which analysts continuously form and revise assumptions in response to evidence (Berret and Munzner, 2024; Grolemund and Wickham, 2014). Analysts rarely begin from blank slates (Kandel et al., 2012; Wongsuphasawat et al., 2019). Instead, they arrive with assumptions about the data (e.g., ”economic growth drives housing demand”) that shape initial inquiry, such as causal hypotheses (e.g., “higher income areas will show greater price stability”), exploratory questions (e.g., “which boroughs deviate from this pattern?”), or domain-driven expectations (e.g., “urban centers tend to recover faster after downturns”(Rule et al., 2018; Heuer, 1999; Fu et al., 2025). In other words, data exploration entails a cycle of interpretive reasoning: analyzing information to extract meaning, questioning or refining these assumptions, and developing evidence-based explanations that guide the course of exploration.
@@
However, this cycle of interpretive reasoning often remains implicit (North et al., 2011; Dou et al., 2009). Analysts may hold assumptions in memory, pursue lines of inquiry without articulating their rationale, or shift interpretations without documenting the changes that led to them (Lipford et al., 2010). When reasoning stays internal, assumptions can harden while still being unverified. Analysts may explore too narrowly, following a single line of inquiry without considering alternatives, or they may selectively seek evidence that confirms initial hunches while overlooking contradictions (Nickerson, 1998; Wall et al., 2017). As exploration deepens, the cognitive burden increases: earlier insights are forgotten, the rationale behind interpretive shifts fades, and promising questions get abandoned without resolution. Without mechanisms to externalize and revisit their evolving thinking, analysts struggle to maintain awareness of how their interpretations have developed, what alternatives they’ve considered, or which questions remain unresolved.
@@
Constructing narrative offers a structure for making this interpretive reasoning explicit. By structuring questions, interpretations, and evidence into narrative form, internal thought becomes tangible statements that can be inspected, revised, and linked to supporting evidence (Murray, 2003; Scaife and Rogers, 1996; Liu and Stasko, 2010). This externalization creates opportunities for reflection, allowing analysts to trace how interpretations evolved, compare competing explanations, and identify gaps in their reasoning (Srinivasan et al., 2018; Islam et al., 2024).
@@
However, most analytic tools do not support analysts’ narrative as a scaffold for interpretive reasoning. They commonly treat narrative as a static output, summarizing finished actions or presenting isolated facts (Keith Norambuena et al., 2022; Wootton et al., 2024). Even with recent advances in AI-assisted data analysis that integrate semantics, natural language, and visualization to support more human-centered exploration (Setlur et al., 2024; Guo et al., 2024; Fu et al., 2025; Setlur, 2025), existing systems remain largely reactive: generating visualizations on demand, assembling analysts’ narratives retrospectively from logs (Lisnic et al., 2025; Wang et al., 2025; Zhao et al., 2021), and responding to explicit prompts rather than participating in meaning construction as it unfolds (Li et al., 2023; Feng et al., 2017; Madanagopal et al., 2019). This keeps exploration (data interaction) separate from interpretation (meaning construction) (Tory et al., 2021; Batch and Elmqvist, 2018), leaving interpretive reasoning invisible and unsupported (Narechania et al., 2025a; Sun et al., 2022). While provenance systems document interaction histories, they capture what analysts do but not why. They lack the narrative logic that connects observations into coherent interpretations (Ragan et al., 2015; North et al., 2011). Without structured externalization, analysts often face difficulties engaging in systematic reflection (Dou et al., 2009; Tory and Moller, 2004), which can lead to premature conclusions or leave assumptions unexamined (Nickerson, 1998; Wall et al., 2017; Li et al., 2025). When reasoning remains implicit, decisions lack the traceability needed for reproducibility and accountability (Gotz and Zhou, 2009; Ragan et al., 2015), which are crucial in fields like journalism (Fu and Stasko, 2023) and intelligence analysis (Saletta et al., 2020). Ultimately, current approaches limit the ability to share, validate, and build upon insights (Willett et al., 2011; Mathisen et al., 2019).
@@
This work addresses these shortcomings by exploring a new interaction paradigm that scaffolds interpretive reasoning as it unfolds. By treating narrative construction as the primary interface of the sensemaking process, our approach makes evolving assumptions explicit, surfaces contradictions, and preserves reasoning paths that are otherwise lost. To achieve this, we draw on insights from theories in narrative psychology (Murray, 2003), external cognition (Scaife and Rogers, 1996; Liu and Stasko, 2010), and constructionist learning (Harel and Papert, 1991; Bartram et al., 2021), which collectively emphasize how externalized narratives can serve both as a cognitive medium for reasoning and as an artifact that preserves its evolution.
@@
Grounded in these perspectives, we propose a new framework—Narrative Scaffolding—that repositions narrative construction as the central interface for data exploration and reasoning. Unlike traditional tools that log low-level interactions or assemble stories retrospectively, Narrative Scaffolding treats interpretation as a dynamic, evolving process that must be externalized and structured to support systematic reasoning. We instantiate this framework in a system that enables analysts to write narrative sentences as first-class analytic actions, with each sentence bidirectionally linked to visual evidence and tracked questions. The system generates semantically-aligned visualizations from natural language, captures insight provenance through a navigable timeline, tracks inquiry evolution through an interactive board, and detects opportunities for branching when reasoning diverges.
@@ 2.1. Narrative in Interpretive Reasoning
Narrative psychology posits that humans are fundamentally wired to make sense of information through constructing coherent narratives to connect events, observations, and meaning (Murray, 2003; Murray et al., 2015). This narrative mode of thought serves not merely as a communication device but as a cognitive tool for organizing experience and building understanding. Sensemaking research supports this view, showing how people iteratively generate and revise interpretations as they build mental models (Pirolli and Card, 2005; Klein et al., 2007b), structuring their understanding around storylines that link goals, observations, and conclusions (Dervin, 1998).
This narrative mode of reasoning manifests across diverse analytical contexts. In journalism, reporters develop story angles that guide their investigation and evidence gathering (Attfield and Dowell, 2003). In scientific inquiry, researchers frame hypotheses narratively, iterating on explanatory models as new data emerges (Padian, 2018; Dahlstrom, 2014). In data analysis, analysts similarly develop and iterate on storylines for organizing complex information and guiding exploratory decisions (Srinivasan et al., 2018; Islam et al., 2024; Fu and Stasko, 2023). Across these domains, narrative functions as an active reasoning scaffold rather than merely a final output format.
Recent empirical studies validate this narrative-driven reasoning in data exploration. Even without explicit tool support, users spontaneously construct storylines during analysis (Srinivasan et al., 2018). Observational research shows analysts following interpretive arcs, revisiting assumptions as evidence emerges (Islam et al., 2024), and beginning with specific narratives that guide their data exploration (Fu et al., 2025). These findings confirm that reasoning unfolds through storyline construction during active exploration, not merely during final communication (Stolper et al., 2014; Zhao et al., 2021).
@@
2.2. Supporting Analytic Reasoning
@@
Supporting analytic reasoning requires capturing not just what analysts do, but why they do it and how their understanding evolves. Provenance tracking addresses this by recording the history of how analytical conclusions were reached, capturing the sequence of actions, decisions, and intermediate states that led to the final result (Heer et al., 2008). In nonlinear, iterative exploration, provenance provides a structured record supporting reproducibility, error recovery, collaboration, and accountability (Ragan et al., 2015; North et al., 2011; Huang and Nobre, 2025; Gotz and Zhou, 2009; Mathisen et al., 2019; Camisetty et al., 2018). However, achieving these goals depends critically on what information can actually be captured.
@@
Ragan et al. (Ragan et al., 2015) proposed a foundational framework categorizing provenance into five types: data, visualization, interaction, insight, and rationale. While the first three can be captured through system instrumentation, the latter two involve internal reasoning that is inherently more challenging to observe (Dou et al., 2009; North, 2006; Willett et al., 2011). Current systems excel at the observable levels. Workflow-based systems like VisTrails (Callahan et al., 2006) model analysis steps in structured pipelines, though they impose rigid workflows less suited to open-ended exploration. Interactive tracking systems (Heer et al., 2008; Cutler et al., 2020; Gratzl et al., 2016) dynamically log actions and states, while recent approaches like ProvenanceWidgets (Narechania et al., 2025b) and ViStruct (Huang and Nobre, 2025) embed tracking directly into interfaces. However, these systems primarily capture observable behaviours while providing limited support for representing underlying reasoning.
@@
To capture insight-level provenance, systems like Calliope (Shi et al., 2020) support explicit argument construction by linking claims to data evidence, while CZSaw (Kadivar et al., 2009) enables analysts to annotate findings and attach rationale to visual states. However, these approaches require users to explicitly articulate insights at specific moments, creating friction between ongoing exploration (Ragan et al., 2015). Insights are often documented as static snapshots instead of representing the dynamic nature of evolving thought processes (North et al., 2011).
@@
This leaves a fundamental gap in capturing interpretive reasoning as it unfolds. Provenance systems excel at documenting what users do but struggle with why they do it. This interpretive evolution (the narrative construction of meaning) is central to how analysts actually reason. When systems cannot capture or support this narrative evolution, reasoning processes remain implicit or fragmented, making it challenging to revisit rationale and trace interpretive evolution (Tory and Moller, 2004; North, 2006; Stolper et al., 2014). Our approach addresses this gap by enabling systems to capture insight-level provenance as reasoning naturally unfolds through writing.
@@ 2.3. LLM-Assisted Data-Driven Sensemaking
Complementary to narrative storyline construction and systematic provenance tracking, mixed-initiative analysis systems address interpretive reasoning through human-AI collaboration (Fu et al., 2025; Lisnic et al., 2025; Latif et al., 2021; Wang et al., 2025). These systems shift from passive tools to collaborative partners that generate views and respond to natural language queries during exploration.
Current approaches can be categorized into three main interaction paradigms. Conversational interfaces enable dialogue-based exploration where users pose questions and refine queries through natural language (Setlur et al., 2016; Gao et al., 2015; Narechania et al., 2020; Obeid and Hoque, 2020; Sultanum and Setlur, 2024). Structured reasoning systems support explicit argument construction through claim-evidence linking and branching logic paths (Shi et al., 2020; Jiang et al., 2023), while text-visualization integration approaches facilitate coherence between textual narratives and visual elements (Sultanum et al., 2021; Zhu-Tian and Xia, 2022; Latif et al., 2021; Srinivasan et al., 2025). A key affordance is that text serves as a more expressive interaction medium, enabling iterative communication between users and systems that goes beyond traditional point-and-click interfaces.
A significant focus has emerged on AI-assisted content generation during analysis. These systems leverage LLMs to automatically produce summaries, story fragments, or visual descriptions based on user actions and analytic traces (Fu et al., 2025; Wang et al., 2025; Sultanum and Srinivasan, 2023; Lisnic et al., 2025; Wang et al., 2019). The effectiveness of these content generation approaches depends on factors such as the sophistication of natural language processing, the quality of AI-generated insights, and the seamless integration of human input with system responses (Li et al., 2024). Our framework shifts from generating narratives reactively to scaffolding them proactively, treating narrative construction as the primary interface for reasoning throughout exploration.
@@ 3. Formative Studies
To understand how interpretive reasoning unfolds during data exploration and where current tools create friction, we conducted a formative study that organized these issues into four interpretive challenges. We surveyed 48 analysts to validate the prevalence of challenges and conducted follow-up interviews with 5 participants to examine how reasoning strategies evolve. Below, we detail the procedure for survey and interview studies and describe the identified challenges that informed our design goals.
@@ 3.1.1. Formative Survey ($N=48$)
We surveyed 48 data practitioners ($F_{1}$$F_{48}$) who frequently engage in data exploration. Participants were recruited through Prolific (Prolific, 2024) and compensated at the rate of \$16 USD/hr. The sample (26 male, 22 female, ages 21–55) included data analysts ($n=16$), managers ($n=7$), healthcare professionals ($n=6$), and sales consultants ($n=3$). This ensured diversity across professional domains, capturing a broad range of reasoning contexts. The survey combined Likert items with semi-structured open-ended prompts, probing how participants initiated exploration, adapted their thinking when encountering unexpected findings, managed analytical complexity, and revisited prior insights. (see supplementary materials for the full protocol)
Responses revealed common behavioural patterns: managing multiple exploration paths through workarounds, externalizing reasoning outside the analytic environment, and restarting from scratch when interpretations diverged. To further investigate how interpretive reasoning unfolds over time and what triggers specific analytical decisions, we conducted follow-up interviews with five participants.

### v2

# Narrative Scaffolding: A Narrative-First Framework for Data-Driven Sensemaking

@@ Abstract.
When exploring data, analysts construct narratives about what the data means by asking questions, generating visualizations, reflecting on patterns, and revising their interpretations as new insights emerge. Yet existing analysis tools treat narrative as an afterthought, breaking the link between reasoning, reflection, and the evolving story from exploration. Consequently, analysts lose the ability to see how their reasoning evolves, making it harder to reflect systematically or build coherent explanations. To address this gap, we propose Narrative Scaffolding, a framework for narrative-driven exploration that positions narrative construction as the primary interface for exploration and reasoning. We implemented this framework in a system that externalizes iterative reasoning through narrative-first entry, semantically aligned view generation, and reflection support via insight provenance and inquiry tracking. In a within-subject study ($N=20$), we demonstrated that narrative scaffolding facilitates broader exploration, deeper reflection, and more defensible narratives. An evaluation with visualization literacy experts ($N=6$) confirmed that the system produced outputs aligned with narrative intent and facilitated intentional exploration.
@@ 1. Introduction
Exploring data is fundamentally a sensemaking process that involves practicing interpretive reasoning throughout: in which analysts continuously form and revise assumptions in response to evidence (berret2024iceberg; grolemund2014cognitive). Analysts rarely begin from blank slates (kandel2012enterprise; wongsuphasawat2019goals). Instead, they arrive with assumptions about the data (e.g., ”economic growth drives housing demand”) that shape initial inquiry, such as causal hypotheses (e.g., ”higher income areas will show greater price stability”), exploratory questions (e.g., ”which boroughs deviate from this pattern?”), or domain-driven expectations (e.g., ”urban centers tend to recover faster after downturns”(rule2018exploration; heuer1999psychology; fu2025dataweaver). In other words, data exploration entails a cycle of interpretive reasoning: analyzing information to extract meaning, questioning or refining these assumptions, and developing evidence-based explanations that guide the course of exploration.
@@
However, this cycle of interpretive reasoning often remains implicit (north2011analytic; dou2009recovering). Analysts may hold assumptions in memory, pursue lines of inquiry without articulating their rationale, or shift interpretations without documenting the changes that led to them (lipford2010helping). When reasoning stays internal, assumptions can harden while still being unverified. Analysts may explore too narrowly, following a single line of inquiry without considering alternatives, or they may selectively seek evidence that confirms initial hunches while overlooking contradictions (nickerson1998confirmation; wall2017warning). As exploration deepens, the cognitive burden increases: earlier insights are forgotten, the rationale behind interpretive shifts fades, and promising questions get abandoned without resolution. Without mechanisms to externalize and revisit their evolving thinking, analysts struggle to maintain awareness of how their interpretations have developed, what alternatives they’ve considered, or which questions remain unresolved.
@@
Constructing narrative offers a structure for making this interpretive reasoning explicit. By structuring questions, interpretations, and evidence into narrative form, internal thought becomes tangible statements that can be inspected, revised, and linked to supporting evidence (murray2003narrative; scaife1996external; liu2010mental). This externalization creates opportunities for reflection, allowing analysts to trace how interpretations evolved, compare competing explanations, and identify gaps in their reasoning (srinivasan2018augmenting; islam2024datanarrative).
@@
However, most analytic tools do not support analysts’ narrative as a scaffold for interpretive reasoning. They commonly treat narrative as a static output, summarizing finished actions or presenting isolated facts (keith2022design; wootton2024charting). Even with recent advances in AI-assisted data analysis that integrate semantics, natural language, and visualization to support more human-centered exploration (cooperative_setlur; guo2024talk2data; fu2025dataweaver; setlur2025sigmod), existing systems remain largely reactive: generating visualizations on demand, assembling analysts’ narratives retrospectively from logs (lisnic2025plume; wang2025jupybara; zhao2021chartstory), and responding to explicit prompts rather than participating in meaning construction as it unfolds (onTheFly; hindsight; madanagopal2019analytic). This keeps exploration (data interaction) separate from interpretation (meaning construction) (tory2021finding; gap_batch), leaving interpretive reasoning invisible and unsupported (designProbe; sun2022erato). While provenance systems document interaction histories, they capture what analysts do but not why. They lack the narrative logic that connects observations into coherent interpretations (ragan2015characterizing; north2011analytic). Without structured externalization, analysts often face difficulties engaging in systematic reflection (dou2009recovering; tory2004human), which can lead to premature conclusions or leave assumptions unexamined (nickerson1998confirmation; wall2017warning; confirmation2025). When reasoning remains implicit, decisions lack the traceability needed for reproducibility and accountability (gotz2009characterizing; ragan2015characterizing), which are crucial in fields like journalism (fu2023more) and intelligence analysis (saletta2020role). Ultimately, current approaches limit the ability to share, validate, and build upon insights (willett2011commentspace; mathisen2019insideinsights).
@@
This work addresses these shortcomings by exploring a new interaction paradigm that scaffolds interpretive reasoning as it unfolds. By treating narrative construction as the primary interface of the sensemaking process, our approach makes evolving assumptions explicit, surfaces contradictions, and preserves reasoning paths that are otherwise lost. To achieve this, we draw on insights from theories in narrative psychology (murray2003narrative), external cognition (scaife1996external; liu2010mental), and constructionist learning (harel1991constructionism; bartram2021untidy), which collectively emphasize how externalized narratives can serve both as a cognitive medium for reasoning and as an artifact that preserves its evolution.
@@
Grounded in these perspectives, we propose Narrative Scaffolding: a framework that repositions narrative construction as the central interface for data exploration and reasoning. Unlike traditional tools that log low-level interactions or assemble stories retrospectively, Narrative Scaffolding treats interpretation as a dynamic, evolving process that must be externalized and structured to support systematic reasoning. We instantiate this framework in a system that enables analysts to write narrative sentences as first-class analytic actions, with each sentence bidirectionally linked to visual evidence and tracked questions. The system generates semantically-aligned visualizations from natural language, captures insight provenance through a navigable timeline, tracks inquiry evolution through an interactive board, and detects opportunities for branching when reasoning diverges.
@@ 2.1. Narrative in Interpretive Reasoning
Narrative psychology posits that humans are fundamentally wired to make sense of information through constructing coherent narratives to connect events, observations, and meaning (murray2003narrative; murray2015narrative). This narrative mode of thought serves not merely as a communication device but as a cognitive tool for organizing experience and building understanding. Sensemaking research supports this view, showing how people iteratively generate and revise interpretations as they build mental models (pirolli2005sensemaking; klein2007dataframe), structuring their understanding around storylines that link goals, observations, and conclusions (dervin1998sense).
This narrative mode of reasoning manifests across diverse analytical contexts. In journalism, reporters develop story angles that guide their investigation and evidence gathering (attfield2003information). In scientific inquiry, researchers frame hypotheses narratively, iterating on explanatory models as new data emerge (padian2018narrative; dahlstrom2014using). In data analysis, analysts similarly develop and iterate on storylines for organizing complex information and guiding exploratory decisions (srinivasan2018augmenting; islam2024datanarrative; fu2023more). Across these domains, narrative functions as an active reasoning scaffold rather than merely a final output format.
Recent empirical studies validate this narrative-driven reasoning in data exploration. Even without explicit tool support, users spontaneously construct storylines during analysis (srinivasan2018augmenting). Observational research shows analysts following interpretive arcs, revisiting assumptions as evidence emerge (islam2024datanarrative), and beginning with specific narratives that guide their data exploration (fu2025dataweaver). These findings confirm that reasoning unfolds through storyline construction during active exploration, not merely during final communication (progressiveInsight; zhao2021chartstory).
@@
2.2. Capturing Analytic Reasoning
@@
Supporting analytic reasoning requires capturing not just what analysts do, but why they do it and how their understanding evolves. Provenance tracking addresses this by recording the history of how analytical conclusions were reached, capturing the sequence of actions, decisions, and intermediate states that led to the final result (heer2008graphical). In nonlinear, iterative exploration, provenance provides a structured record supporting reproducibility, error recovery, collaboration, and accountability (ragan2015characterizing; north2011analytic; huang2025vistruct; gotz2009characterizing; mathisen2019insideinsights; camisetty2018enhancing). However, achieving these goals depends critically on what information can actually be captured.
@@
Ragan et al. (ragan2015characterizing) proposed a foundational framework categorizing provenance into five types: data, visualization, interaction, insight, and rationale. While the first three can be captured through system instrumentation, the latter two involve internal reasoning that is inherently more challenging to observe (dou2009recovering; north2006toward; willett2011commentspace). Current systems excel at the observable levels. Workflow-based systems like VisTrails (callahan2006vistrails) model analysis steps in structured pipelines, though they impose rigid workflows less suited to open-ended exploration. Interactive tracking systems (heer2008graphical; cutler2020trrack; gratzl2016visual) dynamically log actions and states, while recent approaches like ProvenanceWidgets (widgets) and ViStruct (huang2025vistruct) embed tracking directly into interfaces. However, these systems primarily capture observable behaviors while providing limited support for representing underlying reasoning.
@@
To capture insight-level provenance, systems like Calliope (shi2020calliope) support explicit argument construction by linking claims to data evidence, while CZSaw (kadivar2009capturing) enables analysts to annotate findings and attach rationale to visual states. However, these approaches require users to explicitly articulate insights at specific moments, creating friction between ongoing exploration (ragan2015characterizing). Insights are often documented as static snapshots instead of representing the dynamic nature of evolving thought processes (north2011analytic).
@@
This leaves a fundamental gap in capturing interpretive reasoning as it unfolds. Provenance systems excel at documenting what users do but struggle with why they do it. This interpretive evolution (the narrative construction of meaning) is central to how analysts actually reason. When systems cannot capture or support this narrative evolution, reasoning processes remain implicit or fragmented, making it challenging to revisit rationale and trace interpretive evolution (tory2004human; north2006toward; progressiveInsight). Our approach addresses this gap by enabling systems to capture insight-level provenance as reasoning naturally unfolds through writing.
@@ 2.3. LLM-Assisted Data-Driven Sensemaking
Complementary to narrative storyline construction and systematic provenance tracking, mixed-initiative analysis systems address interpretive reasoning through human-AI collaboration (fu2025dataweaver; lisnic2025plume; latif2021kori; wang2025jupybara). These systems shift from passive tools to collaborative partners that generate views and respond to natural language queries during exploration.
Current approaches can be categorized into three main interaction paradigms. Conversational interfaces enable dialogue-based exploration where users pose questions and refine queries through natural language (setlur2016eviza; gao2015datatone; narechania2020nl4dv; obeid2020chart; sultanum2024instruction). Structured reasoning systems support explicit argument construction through claim-evidence linking and branching logic paths (shi2020calliope; jiang2023graphologue), while text-visualization integration approaches facilitate coherence between textual narratives and visual elements (sultanum2021leveraging; zhu2022crossdata; latif2021kori; pluto). A key affordance is that text serves as a more expressive interaction medium, enabling iterative communication between users and systems that goes beyond traditional point-and-click interfaces.
A significant focus has emerged on AI-assisted content generation during analysis. These systems leverage LLMs to automatically produce summaries, story fragments, or visual descriptions based on user actions and analytic traces (fu2025dataweaver; wang2025jupybara; sultanum2023datatales; lisnic2025plume; wang2019datashot). The effectiveness of these content generation approaches depends on factors such as the sophistication of natural language processing, the quality of AI-generated insights, and the seamless integration of human input with system responses (li2024we). Our framework shifts from generating narratives reactively to scaffolding them proactively, treating narrative construction as the primary interface for reasoning throughout exploration.
@@ 3. Formative Studies
To understand how interpretive reasoning unfolds during data exploration and where current tools create friction, we conducted a formative study that organized these issues into four interpretive challenges. We surveyed 48 analysts to validate the prevalence of challenges and conducted follow-up interviews with 5 participants to examine how reasoning strategies evolve. A preliminary version of these findings was reported in (huang2025toward). Below, we detail the procedure for survey and interview studies and describe the identified challenges that informed our design goals.
@@ 3.1.1. Formative Survey ($N=48$)
We surveyed 48 data practitioners ($F_{1}$$F_{48}$) who frequently engage in data exploration. Participants were recruited through Prolific (prolific2024) and compensated at the rate of \$16 USD/hr. The sample (26 male, 22 female, ages 21–55) included data analysts ($n=16$), managers ($n=7$), healthcare professionals ($n=6$), and sales consultants ($n=3$). This ensured diversity across professional domains, capturing a broad range of reasoning contexts. The survey combined Likert items with semi-structured open-ended prompts, probing how participants initiated exploration, adapted their thinking when encountering unexpected findings, managed analytical complexity, and revisited prior insights. (see supplementary materials for the full protocol)
Responses revealed common behavioral patterns: managing multiple exploration paths through workarounds, externalizing reasoning outside the analytic environment, and restarting from scratch when interpretations diverged. To further investigate how interpretive reasoning unfolds over time and what triggers specific analytical decisions, we conducted follow-up interviews with five participants.